Night Time 20140322

6:00 - II test of new collimation tables. JH wants to check the preset acquire accuracy with and without the dichroic in.

Telescope pointed to BS9143, mr=14, 62 elev.

Preset acquire (agw on axis) in parallactic mode, position angle 90deg (PE90). The rotator angle is at 84.8deg: not good, JH prefers having PE0.

6:20 - New preset acquire with PE0 on the same target, rotator angle is at 3.39deg.

Preset active and guiding on.

Dichroic slided in and new preset acquire authorizing LUCI@ARGOS focal station: we find the star within 3": good.

6:35 - dichroic slided out and preset acquire authorizing LUCI@RFBG: we find the star within 1": the test of the new collimation table made by JH is succesful!

6:50 - Handover to Laser operation, spotters outside and shutter test performed.

6:57 - Start of laser propagation, laser pointing adjustment with MAT, cirrus at 10km make laser beams blobby. On PatCams we see M2 spider. TN: 20140322_070531

13444222674_46701c6009_o.jpg patrol_camera_blue.jpeg

Loaded reconstructor for PatCam vs LM1 offload: 20140320_063219, 3 manual iterations.

TN: 20140322_071605 spot are closer as possible to LGSW FoV.

We try to acquire the IM between the PatCam and the LAS periscopes, other adjustment with LM1.

20140322_072523: PatCam to periscope IM with 5arcsec movement, reconstructor loaded.

We try to offset the LGS position error to periscopes.

TN: 20140322_072847 fit is not valid!!!

RED PatCam offset evaluated by hand and periscopes moved by hand: the spot is properly centered on PatCam but we see no light on PnCCD (just the Blue pupil).

TN: 20140322_074331 3 LGS recentered

LAN focus stage set to -100k steps

LM1 moved to recenter just BLUE pupil on PnCCD

Time of flight test: 90us better than 70us

Test of ToF sweep between 9 and 16km: too fast the change of ToF. RED and YELLOW pupils change behaviour with sweep. BLUE one stay fixed. We try also a hand sweep. 90km seems the better position.

We sweep also the LAN focus. We put it to -60k.

TN: 20140322_080643

We close both PI and pupil loops

TN: 20140322_080832

20140322_080832_lgsw_ccdframes_averaged.jpg

TN: 20140322_081238 just pupil loop closed

20140322_081238_lgsw_ccdframes_averaged.jpg

We try to record the IM between PI and periscope: 20140322_082121

The measure offset are too big. We try a IMX0.1

8:38 - periscope gain at 0.9, bad. gain at 0.6. The fit procedure is not working when the spot is in the LGSW FoV. A quad cell algorithm could be implemented.

8:43 - Invalid fit fixed by LB putting 0 to the measured position in case of non converging fit...

8:46 - We try to close the pupil and jitter loops

8:56 - Sweeo ToF: 90us choosed by eye, this corresponds to 13.5 km

8:58 - We try to apply a static shape to M2: 2um iof Z5

TN: 20140322_090131

New slope_null with M2 flat: 20140322_090256

Then we apply 10um of Z5 to M2

9:02 - we try to seep the LAN focus taking snapshots to measure the FHWM:

TN: 20140322_091455 LAN @ -180k steps, not good

TN: 20140322_091527 LAN @ -160k

TN: 20140322_091552 LAN @ -140k

lgsw_ccdframes_averaged_160k-140k.jpg

TN: 20140322_091827 LAN @ -120k

TN: 20140322_091859 LAN @ -100k

TN: 20140322_091936 LAN @ -80k

TN: 20140322_092105 LAN @ -60k, we recenter the LGS, not good

TN: 20140322_092133 LAN @ -60k

TN: 20140322_092209 LAN @ -40k

We recenter the LGS

TN: 20140322_092408 LAN @ -20k

Now we put LGS on PatCams and we sweep again:

TN: 20140322_094639 LAN @ -180k steps, spot on PatCams

TN: 20140322_094734 LAN @ -160k steps, spot on PatCams

TN: 20140322_094757 LAN @ -140k steps, spot on PatCams

TN: 20140322_094839 LAN @ -120k steps, spot on PatCams

TN: 20140322_094931 LAN @ -100k steps, spot on PatCams

TN: 20140322_095009 LAN @ -80k steps, spot on PatCams

TN: 20140322_095055 LAN @ -60k steps, spot on PatCams

20140322_094639_focusStagePositionVsPatCamSpotSize.png

Yellow LGS appears a little more out of focus with respect the other 2 LGS...

The roundish spot is at -120k.

Recenter LGS with LM1 and periscope loops.

TN: 20140322_102601 mix of internal source + LGS, we're vignetting the red LGS

11:39 - Switched on FLAO, telescope is collimated on WT10_292 (mr=6.5, 14.5deg from Zenith), AO loop closed and new flat saved for M2 (ARGOS and ARGOS1)

New slope_null with flattened M2: 20140322_120511

TN: 20140322_120026 snapshot with FLAO on and Jitter Open Loop

12:31 - Telescope back to Zenith having set ARGOS1 flat, M2 is back to banana shape...

TN: 20140322_125803 good snapshot to define subapertures.
Topic revision: r3 - 21 Feb 2015, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback